i need your help the question is the foundation of a building is in the shape of two adjacent squares. the perimeter of the foundation is 80 m and its area is 325 m^2. find the length of a side of each square. these are the equations i have 4x+ 2y= 80 and x^2 + y^2= 325
Answer by Alan3354(69443)

the foundation of a building is in the shape of two adjacent squares. the perimeter of the foundation is 80 m and its area is 325 m^2. find the length of a side of each square. these are the equations i have 4x+ 2y= 80 and x^2 + y^2= 325
-----------
4x + 2y = 80 is good for the perimeteter, but
x^2 + y^2 = 325 is a circle.
-----
2xy = 325 is the Area